OpenStudy (imtiaz7):

A body of weight 200 newton falls from the height of 15 m and penetrate 2 cm into sand. What average force is exerted on it by the sand, neglect air fraction.

OpenStudy (michele_laino):

here, the work done by the average force \(R\), has to be equal to the potential energy of the falling object, so we can write this equation: \[\Large mgh = Rd\] where \(d=2\) cm= \(0.02\) meters. So, we can write this: \[\Large R = \frac{{mgh}}{d} = \frac{{Ph}}{d} = \frac{{2000 \cdot 15}}{{0.02}} = ...?\]